OpenStudy (anonymous):

help graphing quadradtic function y+X^2+5x+2=0

OpenStudy (anonymous):

the vertex of the function is:\[\frac{ -b }{ 2a }\]

OpenStudy (anonymous):

so if you write your equation in terms of y you get y = -x^2-5x-2

OpenStudy (anonymous):

so what is the a term?

OpenStudy (anonymous):

can u tell me?

OpenStudy (anonymous):

-x

OpenStudy (anonymous):

the x-coordinate of the vertex is -b/2a so we need to know the a nd be terms

OpenStudy (anonymous):

yah so a is -1 right?

OpenStudy (anonymous):

what is b?

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

yup so the x-cooridnate of the vertex of the parabola is -b/2a which is?

OpenStudy (anonymous):

-(-5)/2(-1)=?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

2.5

OpenStudy (anonymous):

actually it is -2.5

OpenStudy (anonymous):

cuz its 5/2(-1) = -2.5

OpenStudy (anonymous):

how

OpenStudy (anonymous):

-(-5)/2(-1)

OpenStudy (anonymous):

-(-5) is 5

OpenStudy (anonymous):

2(-1) is -2

OpenStudy (anonymous):

so it is 5/-2 which is -2.5

OpenStudy (anonymous):

ok so the x-coordinate of the vertex is -2.5 so how do u think u would find the y-coordinate?

OpenStudy (anonymous):

ok i get it

OpenStudy (anonymous):

plug in -2.5 for x

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so what do you get?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

i got 4.25

OpenStudy (anonymous):

-(2.5)^2-5(2.5)-2

OpenStudy (anonymous):

yah wolfram says its 4.25 too: http://www.wolframalpha.com/input/?i=vertex+of+parabola+y%3D-x%5E2-5x-2

OpenStudy (anonymous):

tok i have to go

OpenStudy (anonymous):

so after you find the vertex find the y-intercepts where x=0

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so the vertex is (-2.5,4.25) so graph that

OpenStudy (anonymous):

so plug in x=0 to find at which y-coordinate it crosses the yaxis

OpenStudy (anonymous):

so can u tell me where that is?

OpenStudy (anonymous):

its -2 right?

OpenStudy (anonymous):

if you plug in x=0 you get -2

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so it crosses the y-axis at x=-2

OpenStudy (anonymous):

i mean at y= -2

OpenStudy (anonymous):

so graph that

OpenStudy (anonymous):

it would be 2 below on the yaxis

OpenStudy (anonymous):

i have to plot 3 other points

OpenStudy (anonymous):

ok find the x-intercept where y=0

OpenStudy (anonymous):

didnt learn that

OpenStudy (anonymous):

just plug in some number for x and find the y-coordinate

OpenStudy (anonymous):

and that will be a point on the parabola

OpenStudy (anonymous):

then do that for 2 other points

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then you just connect them with curves and then you are done

OpenStudy (anonymous):

do you get it now?

OpenStudy (anonymous):

YES